OpenStudy (anonymous):

Can anyone help me solve this it is a linear equation I had to Create a linear equation in two variables in the form y = mx + b to graph using the slope intercept form and now I have to solve it and I am not sure how to do it?PLEASE HELP ME SOLVE THIS ---->2(2x +5) -2(x-10)=40

OpenStudy (anonymous):

multiply out and it comes to 4x+10-2x+20=40 can you atleast solve from here?

OpenStudy (anonymous):

first multiply out, then combine like terms \[2(2x+5)-2(x-10)=40\] \[4x+10-2x+20=40\] \[2x+30=40\] \[2x=10\] \[x=5\] this is not a line, it is a number

OpenStudy (anonymous):

so 2x =10 x = 5 yea this is a number, not a line